OpenStudy (anonymous):

What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm. is it arc measure/360 = circumference or area??

OpenStudy (anonymous):

L = r* theta where theta = arc measure theta = L / r = 4.189 / 3 radians multiply this by 180/pi gives the arc measure in degrees this comes to 80 degrres D

OpenStudy (anonymous):

You have arclength and radius, and you can use 3.14 for pi. At this point, it's just simple arithmetic.

OpenStudy (anonymous):

Now we use this formula s = (t/360)*C where.. s = arc length t = central angle C = circumference of entire circle

OpenStudy (anonymous):

That's the unknown you're solving for. It turns out that when you say "measure of an arc" you also mean "the measure of the central angle"

OpenStudy (anonymous):

s = (t/360)*C s = (t/360)*18.84 ... plug in the circumference you just found 4.189 = (t/360)*18.84 ... plug in the arc length given now solve for t